# Super-intelligence or Superstition? Exploring Psychological Factors Underlying Unwarranted Belief in AI Predictions

### v2

@@ 2 Results
On the other hand, familiarity with the prediction sources (AI, astrology, personality psychology), which measures the level of self-reported familiarity/prior knowledge in each source, did not have a significant effect on perceived validity (-0.02, 95% CI [-0.11, 0.07], p = 0.650) across all sources. The familiarity scale ranged from 1 to 5 (Mean = 2.96, SD = 1.05).
The interaction between familiarity and subscales shows that perceived personalization increases slightly but significantly by 0.05 points (95% CI [0.00, 0.11], p = 0.035), while differences were not statistically significant for perceived reliability (0.02, 95% CI [-0.06, 0.10], p = 0.685) and perceived usefulness (0.01, 95% CI [-0.06, 0.08], p = 0.781).
@@ 2.9 Other results
Gullibility was not found to be a significant predictor of perceived validity of AI predictions (-0.02, 95% CI [-0.04, 0.01], p = 0.128). While the interaction between gullibility and Personalization subscale was significant (0.01, 95% CI [0.00, 0.02], p = 0.021), the results were inconclusive when compared to the main effects. Interactions between gullibility and prophecy source were not significant (Astrology: -0.00, 95% CI [-0.02, 0.02], p = 0.977, Personality: -0.02, 95% CI [-0.04, 0.00], p = 0.096), suggesting the effects were similarly insignificant across different prophecy sources. Gullibility scale ranged from 6 to 39 (Mean = 14.68, SD = 7.69).
@@
For gender, we did not find significant main effects relative to the Female reference level. While male participants were associated with lower perceived validity scores for AI predictions than female participants, it was not statistically significant (-0.31, 95% CI [-0.69, 0.06], p = 0.103). However, they tend to perceive them as more reliable than female participants (0.41, 95% CI [0.14, 0.68], p = 0.003). Interactions with other subscales were not significant, suggesting they were similar to the perceived validity baseline. Interactions between gender and prophecy source were not significant.
There were no significant main effects for level of education compared to the Bachelor reference level. Interactions between education level and subscale showed some significant effects between High school and Personalization (-0.23, 95% CI [-0.40, -0.06], p = 0.009), and Less than high school and Reliability (-0.89, 95% CI [-1.56, -0.22], p = 0.009), but these results did not lead to conclusive results.
@@ 3 Discussion
Belief is subjective and context-dependent; it is also socially constructed through external stimuli such as media narratives, contextual factors, and internal disposition [48, 49]. When looking at belief, trust, and reliance in AI predictions, it is important to consider not just the features of the AI system, but also the characteristics of the users and the broader context. The results from our study show that people who are more likely to believe in astrology- and personality-based predictions were more likely to believe in AI predictions. Moreover, we found that belief in AI predictions on personal behavior can be influenced by various psychological, cognitive, and contextual factors that also drive belief in astrology- or personality-based predictions. By breaking down the concept of believability into four subscales (perceived validity, reliability, usefulness, and personalization), which we further elaborate in Section 7.3.1, we were able to observe interesting patterns and interactions with different subscales.
Cognitive style. While prior literature seemed to suggest that an analytic cognitive style would lead to more skepticism in predictions based on astrology and possibly personality [31, 33, 35], our results did not show significant evidence to support these claims. Moreover, cognitive style was not a significant predictor in belief in AI predictions. Thus, our hypothesis (H2) that people with more analytic cognitive style would be more likely to believe in AI predictions than predictions based on astrology and personality was not supported by our findings. This suggests that analytic cognitive style may not necessarily lead to rational skepticism of fictitious predictions, and perhaps there may be a missing link or mechanism between cognitive style and belief/trust in AI predictions, which could be a topic of future study.
Paranormal beliefs. Paranormal beliefs were found to be an influential predictor of belief in predictions across all sources, including AI predictions. While the association may seem unlikely at first, it can be interpreted within the context of the influence of popular AI narratives. The popular portrayal of AI as rational and objective creates the impression that people would also treat AI in a rational way, and thus more scientifically-inclined people would find AI predictions more believable. However, our results show that belief in AI predictions is more closely associated with paranormal beliefs than one’s cognitive style. This result points to the existence of the phenomenon of "rational superstition" in AI that was described in the introduction. Moreover, while the effect of paranormal beliefs on belief in AI predictions was significant across all subscales, the effect was larger for perceived usefulness and reliability. Social science research suggests that the rise in belief in astrology is an indicator of rising uncertainties and anxieties as a result of disintegration of community in the modern era [50]. As such, people turn to pseudo-scientific approaches such as astrology for insights to help them navigate their lives. Our findings suggest that people may perceive AI predictions in a way that is similar to the way they perceive astrology-based predictions, which could help explain the results.
@@
AI Attitude/Trust in AI. Our results showed that people with positive attitudes towards AI found AI predictions more valid, reliable, useful, and personalized. Among the subscales, perceived reliability was found to be most closely related to attitude towards AI. This supports findings from prior literature that positive attitudes toward AI lead to higher reliance in AI predictions [51, 46].
@@
Personality. Our findings using the Big Five personality traits [52] show that conscientiousness had a negative influence on perception of validity, personalization, reliability, and usefulness, while other traits did not have significant effects. This contrasts previous findings by Riedl [30], who found that agreeableness, openness, and extraversion positively influenced trust and high neuroticism negatively impacted trust, while results on conscientious were mixed.
@@
Conscientiousness is defined as “socially prescribed impulse control that facilitates task- and goal-directed behavior, such as thinking before acting, delaying gratification, following norms and rules and planning, organizing, and prioritizing tasks” [53, p. 120]. Prior studies have found a positive relationship between conscientiousness and cognitive ability [54, 55]. Contrary to our expectations that these two traits may mirror each other in direction, it was not the case—cognitive style did not have significant effect on believability in our study, while conscientiousness had a significant negative influence on perception of validity, personalization, reliability, and usefulness.
@@
Interest in the topic of prediction. Our findings show that people’s interest in the topic of prediction is a strong predictor when it comes to their belief in predictions, regardless of the source. While our results do not confirm any causal relationship, it could suggest that the more interested one is in a topic, the more exposure they may have to information about it, and the more likely they may be susceptible to cognitive biases that could influence their belief. Some biases that could explain this tendency are confirmation bias and belief bias. Confirmation bias is when individuals seek out and remember information that confirm their existing beliefs [56]. Belief bias occurs when individuals judge the strength of an argument based on the believability of its conclusion rather than the logical structure of the argument itself [57, 58].
@@
Familiarity/Prior knowledge. Familiarity with the prediction source was not found to be a significant predictor of belief in predictions. Earlier, we discussed how people’s positive AI attitudes and interest in the topic of the prediction were positively correlated with belief in the predictions, which may be explained by its likelihood to fuel certain cognitive biases such as belief bias. On the other hand, an individual’s level of familiarity in AI, astrology, and personality psychology may not be necessarily indicative of their attitude towards the field. For instance, someone who is not acquainted with the details of latest AI developments may have either a utopian or dystopian view of AI’s impact on society, as mentioned in the introduction. Similarly, just because an individual is very familiar with astrology does not inherently mean they are more likely to believe or adopt the predictions, as that may depend on other personal and contextual factors. This suggests two things: 1) simply knowing more about the prediction source does not predict whether one finds a prediction more or less believable, and 2) it does not make one better at calibrating one’s trust, providing implications for the design of trustworthy AI systems.
@@ 4 General Discussion
Our results highlight the irrational side of how humans perceive and believe in AI predictions by comparing the perception of fictitious predictions based on AI, astrology, and personality. Our analysis showed that belief in AI prediction about personal behavior may be more related to paranormal beliefs than cognitive style, and that positive attitudes toward AI and interest in the prediction topic enhance this belief, regardless of familiarity with the sources.
These findings suggest that highly accurate and reliable performance is not a prerequisite for people to put high trust in predictions; models do not need to be completely trustworthy for users to trust them. This disconnect between perception and actual performance raises important discussions about how people perceive and engage with AI systems.
1) Incorrect mental models of AI can override perception of validity.
Belief in AI predictions is often rooted in mental models rather than actual performance. Hoff and Bashir’s three-layer framework of trust in automation—dispositional, situational, and learned—illustrates that system performance contributes to learned trust through active engagement with the system [63]. Our study shows that even without validating system performance, people rated fictitious AI predictions as highly valid, reliable, useful, and personalized. This placebo effect of AI has been observed by other researchers, where mere priming with AI descriptions alters perception and behavior [64, 65, 66, 67].
This underscores the significant role mental models play in shaping trust in AI [23, 68, 69, 70, 71, 72]. Mental models, comprising beliefs and expectations about AI, are influenced by prior beliefs and ongoing interactions [23, 68]. By framing the scenario in certain ways, including the task domain and the information provided about the AI system, we observed that people may perceive fictitious AI predictions to be accurate, personalized, and trustworthy [23].
Moreover, our results suggest that people might view AI predictions similarly to astrology, supporting the idea of "rational superstition" in AI and automation [17]. We argue that this phenomenon potentially stems from the idealized portrayal of AI as enchanted yet rational and deterministic [73], which can bias interactions and reinforce existing beliefs about AI.
The dual-process theory in psychology provides insight into how larger narratives can bias perceptions of AI. This theory describes two types of thinking: fast and intuitive (System 1) and slow and reflective (System 2) [20, 21]. Mental models, shaped by public narratives, may act as heuristics that trigger System 1 thinking, leading to speculation rather than informed reasoning [22]. Incomplete or incorrect mental models, shaped by speculative claims rather than evidence, can lead to overreliance or miscalibrated trust, affecting effective AI use [18, 19, 23, 43].
This highlights the need to demystify AI in popular narratives and promote awareness of both its capabilities and limitations to foster accurate mental models. This leads to the importance of AI literacy efforts and how narratives are shaped by providers, media, etc. As technology advances, users’ mental models should evolve with a balanced level of skepticism to prevent overreliance or undue skepticism.
2) Mental models can influence and be influenced by cognitive biases.
@@
Cognitive biases help explain the irrational belief in AI predictions. Belief bias, where individuals evaluate statements based on prior beliefs rather than logical validity [57, 58], is largely driven by System 1 thinking (especially under time pressure [74]) but can also involve System 2, reinforcing mental models aligned with existing beliefs [75, 21]. This selective processing model explains why people may fail to engage logical thinking when a statement aligns with their beliefs, especially concerning self-related beliefs [76].
@@
The Barnum effect, where individuals believe vague, general statements to be highly accurate for them personally, also plays a role [77]. This effect, studied in contexts like astrology and personality tests [78, 79], can explain how people view generic AI-generated predictions as personalized and valid. Cognitive biases like belief bias and the Barnum effect shape mental models and influence AI reliance [69], while incorrect mental models can, in turn, induce these biases.
3) Design of AI systems must consider mental models and cognitive biases to support the development of an appropriate level of trust.
AI systems should be designed to help users form accurate mental models. In contexts like mental health chatbots, incorrect mental models can lead to unrealistic expectations and greater disappointment among users. Thus, these systems should include mechanisms that promote an appropriate level of trust to avoid over or under-reliance, which should be accompanied by efforts to increase AI literacy [46]. Fortunately, mental models are dynamic and can evolve with more interaction [23, 80].
Awareness of the potential impact of cognitive biases on human-AI interaction is crucial to reduce the risk of negatively exploiting these human tendencies. AI systems that merely validate users’ pre-existing beliefs may be favored but may not support better decision-making, particularly in critical areas like personal finance or health. Thus, we could think about how to design systems that can present information in a way that helps people think critically.
There are also important considerations for designing explainable AI systems. While explanations are intended to build trust, they can inadvertently increase over-trust, especially if users have inaccurate mental models [81]. Studies have shown that belief bias can lead to overreliance on AI explanations when they align with prior beliefs [82]. To mitigate this, explanations should be designed to engage users in critical thinking, perhaps by questioning the user or presenting examples that clarify the AI model’s reasoning [83, 84].
@@
4) Our understanding of human-AI interaction needs to include psychological and contextual factors.
@@ 5 Limitations and Future Work
While our study has presented valuable insights into the phenomenon of "irrational superstition" in AI predictions and cognitive biases, there are several important limitations that can be addressed in future work. First, our experiment was specifically designed to provide predictions on personal behavior. There are some limitations to the extent to which our results can be generalized to the general use of AI. Therefore, future studies could look at how people perceive AI predictions in different contexts and prediction topics, while also exploring the relationship with underlying psychological and cognitive factors at play. For instance, the effect can be tested within the context of AI-assisted decision-making in personal health or well-being, or in educational settings using an AI tutor. We believe that our study opens up more questions about understanding the intricate ways in which humans perceive, trust, and interact with AI systems. Second, like most survey designs, our survey-based approach to measuring psychological and cognitive traits may be affected by social desirability bias (the tendency to respond in a way that seems socially favorable) [85], as well as common method biases [86], and reliance on self-reported data. Therefore, we hope future studies can validate our findings using different methods and approaches. Lastly, while our study mainly focused on user-related factors, future studies could also look at other contextual factors such as social and organizational environments and their impact on user perception and trust.

### v3

@@ 2 Results
On the other hand, familiarity with the prediction sources (AI, astrology, personality psychology), which measures the level of self-reported familiarity/prior knowledge in each source, did not have a significant effect on perceived validity (Estimate = -0.02, 95% CI [-0.11, 0.07], p = 0.650) across all sources. The familiarity scale ranged from 1 to 5 (Mean = 2.96, SD = 1.05).
The interaction between familiarity and subscales shows that perceived personalization increases slightly but significantly by 0.05 points (95% CI [0.00, 0.11], p = 0.035), while differences were not statistically significant for perceived reliability (Estimate = 0.02, 95% CI [-0.06, 0.10], p = 0.685) and perceived usefulness (Estimate = 0.01, 95% CI [-0.06, 0.08], p = 0.781).
@@ 2.9 Other results
Gullibility was not found to be a significant predictor of perceived validity of AI predictions (Estimate = -0.02, 95% CI [-0.04, 0.01], p = 0.128). While the interaction between gullibility and Personalization subscale was significant (Estimate = 0.01, 95% CI [0.00, 0.02], p = 0.021), the results were inconclusive when compared to the main effects. Interactions between gullibility and prophecy source were not significant (Astrology: Estimate = -0.00, 95% CI [-0.02, 0.02], p = 0.977, Personality: Estimate = -0.02, 95% CI [-0.04, 0.00], p = 0.096), suggesting the effects were similarly insignificant across different prophecy sources. Gullibility scale ranged from 6 to 39 (Mean = 14.68, SD = 7.69).
@@
For gender, we did not find significant main effects relative to the Female reference level. While male participants were associated with lower perceived validity scores for AI predictions than female participants, it was not statistically significant (Estimate = -0.31, 95% CI [-0.69, 0.06], p = 0.103). However, they tend to perceive them as more reliable than female participants (Estimate = 0.41, 95% CI [0.14, 0.68], p = 0.003). Interactions with other subscales were not significant, suggesting they were similar to the perceived validity baseline. Interactions between gender and prophecy source were not significant.
There were no significant main effects for level of education compared to the Bachelor reference level. Interactions between education level and subscale showed some significant effects between High school and Personalization (Estimate = -0.23, 95% CI [-0.40, -0.06], p = 0.009), and Less than high school and Reliability (Estimate = -0.89, 95% CI [-1.56, -0.22], p = 0.009), but these results did not lead to conclusive results.
@@ 3 Discussion
Belief is subjective and context-dependent; it is also socially constructed through external stimuli such as media narratives, contextual factors, and internal disposition [49, 50]. When looking at belief, trust, and reliance in AI predictions, it is important to consider not just the features of the AI system, but also the characteristics of the users and the broader context. The results from our study show that people who are more likely to believe in astrology- and personality-based predictions were more likely to believe in AI predictions. Moreover, we found that believability of AI predictions on personal behavior can be influenced by various psychological and contextual factors that also drive belief in astrology- or personality-based predictions.
We conceptualized believability using four subscales—perceived validity, reliability, usefulness, and personalization—each grounded in theoretical and empirical research on persuasion, technology acceptance, and user attitudes (see Section 7.3.1 for details). This conceptualization provides a more comprehensive framework for understanding how users evaluate the credibility of predictions, offering insight into the psychological and contextual factors that influence decision-making.
Cognitive style. Our findings challenge the assumption that an analytic cognitive style leads to greater skepticism toward predictions based on astrology or personality assessments and a preference for AI-generated predictions. Contrary to prior research [33, 35, 31] and our hypothesis (H2), cognitive style was not a significant predictor of belief in any type of prediction, including those from AI. Moreover, we observed no inverse relationship between belief in AI and astrology predictions, suggesting that analytic thinkers do not inherently dismiss pseudoscientific forecasts or display heightened trust in AI predictions. This highlights a gap between cognitive style and rational skepticism, pointing to a potential missing mechanism influencing belief in AI predictions. Interestingly, the significant interaction between cognitive scores and perceptions of reliability and usefulness nearly reversed the direction of the nonsignificant main effect (reference level validity), indicating that the impact of cognitive style may depend on context-specific factors, such as how predictions are evaluated for their reliability and usefulness, rather than reflecting a uniform influence across all prediction types.
Paranormal beliefs. Paranormal beliefs emerged as a strong predictor of belief in predictions across all sources, including AI. Although this association may initially seem unexpected, it can be understood within the context of popular narratives about AI. Despite AI being portrayed as rational and objective, which might lead to the assumption that scientifically inclined individuals would find AI predictions more credible, our findings indicate otherwise. Instead, belief in AI predictions was more closely linked to paranormal beliefs than to cognitive style. This aligns with the concept of "rational superstition" in AI, as discussed in the introduction. Furthermore, the influence of paranormal beliefs was significant across all subscales of AI predictions but was especially pronounced for perceived usefulness and reliability. Social science literature suggests that the growing interest in astrology reflects increasing uncertainties and anxieties stemming from the breakdown of traditional communities [51]. People may turn to pseudoscientific approaches like astrology for guidance amid these uncertainties. Our findings indicate that individuals might view AI predictions similarly to astrology-based predictions in the context of personal behavior forecasting, offering a potential explanation for the observed results.
@@
Attitude toward AI. Our results showed that people with positive attitudes towards AI found AI predictions more valid, reliable, useful, and personalized. Among the subscales, perceived reliability was found to be most closely related to attitude towards AI. This supports findings from prior literature that positive attitudes toward AI lead to higher reliance in AI predictions [52, 47].
@@
Personality. Our findings using the Big Five personality traits [53] show that conscientiousness had a negative influence on perception of validity, personalization, reliability, and usefulness, while other traits did not have significant effects. This contrasts with prior literature on personality’s influence on trust in AI systems that show that agreeableness, openness, and extraversion generally have a positive influence on trust, while high neuroticism tends to negatively affect trust [30]. Results on conscientiousness have been inconclusive, with some studies reporting a positive association [54, 55, 56], while others indicate a negative relationship [57, 58].
@@
Conscientiousness is defined as “socially prescribed impulse control that facilitates task- and goal-directed behavior, such as thinking before acting, delaying gratification, following norms and rules and planning, organizing, and prioritizing tasks” [59, p. 120].
While recent studies suggest a positive relationship between conscientiousness and cognitive ability [60, 61], our results did not Contrary to our expectations that cognitive style and conscientiousness may mirror each other in direction, it was not the case—cognitive style did not have significant effect on believability in our study, while conscientiousness had a significant negative influence on perception of validity, personalization, reliability, and usefulness.
@@
Interest in the topic of prediction. Our findings show that people’s interest in the topic of prediction is a strong predictor when it comes to their belief in predictions, regardless of the source. A potential explanation is that the more interest an individual has about the topic, the more exposure they may have to information about it, and the more likely they may be susceptible to cognitive biases that could influence their belief. Some biases that could explain this tendency are confirmation bias and belief bias. Confirmation bias is when individuals seek out and remember information that confirm their existing beliefs [62]. Belief bias occurs when individuals judge the strength of an argument based on the believability of its conclusion rather than the logical structure of the argument itself [63, 64].
@@
Familiarity/Prior knowledge. Familiarity with the prediction source was not found to be a significant predictor of belief in predictions. Earlier, we discussed how people’s positive AI attitudes and interest in the topic of the prediction were positively correlated with belief in the predictions, which may be explained by its likelihood to fuel certain cognitive biases such as belief bias. On the other hand, an individual’s level of familiarity in AI, astrology, and personality psychology may not be necessarily indicative of their attitude towards the field. For instance, someone who is not acquainted with the details of the latest AI developments may have either a utopian or dystopian view of AI’s impact on society, as mentioned in the introduction. Similarly, just because an individual is familiar with astrology does not inherently mean that they are more likely to believe or adopt the predictions, as it may be influenced by other personal and contextual factors. This suggests two things: 1) simply knowing more about the prediction source does not predict whether one finds a prediction more or less believable, and 2) it does not make one better at calibrating one’s trust, providing implications for the design of trustworthy AI systems.
@@ 4 General Discussion
Our results highlight the irrational ways humans perceive and trust AI predictions, showing that belief in AI predictions may be more influenced by paranormal beliefs and positive attitudes toward AI than by cognitive style or familiarity with the technology. This suggests that highly accurate performance is not necessary for users to trust AI systems, pointing to a disconnect between perception and actual capability. This section discusses practical implications of the results.
1) Incomplete mental models of AI can skew perception of validity, reliability, usefulness, and personalization.
Our findings demonstrate that people’s belief in AI predictions is often shaped by their mental models rather than the actual performance of the AI system. Even with entirely fictitious AI predictions, participants rated them as valid, reliable, useful, and personalized. Building on previous research in the placebo effect of AI, this reflects the significant influence of framing and presentation: users may trust AI based on how it is described rather than its true capabilities [69, 70, 71, 72, 73, 74]. Popular narratives that depict AI as omniscient or highly rational contribute to unrealistic mental models, undermine critical thinking and agency when interacting with the predictions [2, 75].
This finding has practical implications for both AI design and public policy. AI systems should be designed to help users develop accurate mental models by using clear, objective language and offering transparency about the system’s limitations. Overhyped marketing or media portrayal can exacerbate misunderstandings about the actual capabilities of AI systems, leading to misplaced trust. Thus, it is important to support AI literacy initiatives that continually educate the public on realistic AI capabilities and limitations, as well as increase efforts to reduce misleading claims. As technology evolves, users’ mental models must also adapt, incorporating a healthy level of skepticism to foster informed, rational interactions [76].
2) Cognitive biases can reinforce flawed mental models of AI systems.
@@
Cognitive biases provide valuable insight into understanding irrational beliefs about AI. Belief bias is one such cognitive bias, where individuals evaluate information based on prior beliefs rather than logical validity [63, 64]. While primarily driven by fast, intuitive System 1 thinking, particularly under time pressure [77], belief bias can also involve slower, more reflective System 2 processes, which may rationalize or reinforce mental models aligned with existing beliefs [78, 10]. Known as the selective processing model, this helps explain why people may fail to engage logical thinking when a statement aligns with their beliefs, especially concerning self-related beliefs [79].
@@
The Barnum effect, where individuals believe vague, general statements to be highly accurate for them personally, also contributes to this phenomenon [80]. Commonly observed in astrology and personality assessments [81, 82], the Barnum effect may explain why users perceive generic AI-generated predictions as personalized and valid. These cognitive biases, including belief bias and the Barnum effect, shape and reinforce flawed mental models of AI, leading to overreliance on its outputs [83, 13, 11, 44, 84]. In turn, these incorrect mental models can reinforce cognitive biases, creating a feedback loop that distorts rational engagement with AI.
These cognitive biases highlight why transparency in AI is not enough. While interpretability and explainability efforts aim to make AI systems clearer, explanations that simply confirm users’ existing biases can amplify over-trust rather than promote understanding [85, 86]. Effective explanations could go beyond clarifying AI processes to encourage thoughtful, critical engagement with AI outputs [87].
3) Promoting critical engagement with AI.
As our reliance on AI-assisted decision-making grows, our capacity for independent critical thinking may be compromised [88]. To counteract this, AI systems must be designed with an awareness of human mental models and cognitive biases to support the development of an appropriate level of trust and foster critical engagement.
Recent research in human-AI collaboration propose promising directions. For instance, interactive feedback mechanisms that offer counterfactual scenarios [89] or pose follow-up questions could prompt users to reconsider their assumptions. Deliberation nudges, such as a "cool-off" period for high-stakes decisions, could also slow down quick judgments and encourage more thoughtful engagement [90]. Additionally, "seamful" design approaches that strategically exposes mismatches and mistakes of AI "to support user agency, re-configuration, and appropriation" [91]. Moreover, presenting uncertainty or confidence levels with AI predictions can encourage users to critically evaluate the outputs rather than blindly accept them [92]. Robust evaluation frameworks and benchmarks should be developed to validate the effectiveness of these approaches in real world systems.
From a policy standpoint, advancing AI literacy is essential to equip the public with the tools to critically engage with AI systems. Initiatives should include clear, accessible guidelines that foster public awareness of the implications of AI-augmented cognition and encourage informed discourse.
The notion that humans are inherently irrational while machines are rational is a flawed perspective. Both humans and AI can exhibit rational and irrational behaviors, and human-machine interactions are shaped by a complex interplay of these tendencies. The goal of AI design should be to support informed and critical decision-making, helping users navigate these complexities in ways that enhance well-being and minimize risks.
@@
4) Our understanding of human-AI interaction should include psychological and contextual factors.
@@
Future research should prioritize interdisciplinary collaboration to delve deeper into the psychological and cultural dimensions of human-AI interactions. Psychologists and cognitive scientists can identify cognitive biases that shape user behavior, while sociologists and anthropologists can explore the influence of societal narratives on public perceptions of AI. Policymakers should actively support such research to better understand the broader societal implications of human-AI interactions. Additionally, AI system designers and developers must integrate these insights to create systems that align with realistic user expectations, fostering trust and effective collaboration between humans and AI.
@@ 5 Limitations and Future Work
First, our experiment focused on predictions about personal behavior to facilitate comparisons with astrology and personality-based predictions. This specific design may limit generalizability. Future research should explore perceptions of AI predictions in diverse contexts and domains, such as health, education, and workplace decision-making, as well as across different interaction settings (e.g., implicit versus explicit predictions). Examining psychological and cognitive factors in these areas could yield richer insights into belief formation and trust dynamics.
Second, our findings highlight the need for research into how cognitive style interacts with context-specific factors, such as perceptions of reliability and usefulness, in shaping beliefs about predictions. The significant interaction observed, which nearly reversed the nonsignificant main effect of validity, suggests that analytic thinkers may evaluate predictions flexibly rather than uniformly. Future studies could explore whether analytic cognitive styles involve distinct evaluation strategies for different prediction sources or whether contextual features, like framing or perceived expertise, influence these judgments. Understanding these mechanisms could inform interventions aimed at enhancing critical thinking and improving discernment between credible and pseudoscientific claims.
Third, our study primarily investigated user-related factors, leaving external influences such as social and organizational environments relatively unexplored. Future research could investigate how these broader contexts shape user perceptions, engagement, and trust in AI predictions, particularly in collaborative or high-stakes decision-making scenarios.
Finally, like most survey-based studies, ours may be affected by biases inherent to self-reported data, including social desirability bias [94] and common method bias [95]. To address these limitations, future studies could incorporate alternative methodologies, such as behavioral experiments or longitudinal designs, to validate and extend our findings.